Time period of a simple pendulum of length L is T1 and time period of a uniform rod of the same length is pivoted at one end an oscillating in a vertical plane is T2 . Amplitude of oscillation in both the cases is small. Then T1/T2 is
A) 1/3½        B) 1
C) (4/3)½      D)  (3/2)½

Grade:11

1 Answers

Arun
5 years ago
For T1 we have to use formula 

T 1 = √L/g as it is a simple pendulum.
T 2 = √I/mgy as it is compound pendulum.

T 1 = √L/g

T 2 = √mL^2/3mgL
= √2L/3

T 1 / T 2 = √L/g ÷ √2L/3

= √3/2.